OpenStudy (dannimarie):

A mine extracts 2 tons of ore in an hour. Of the total ore extracted in a day, 4 tons is purchased by the government and more than 10 tons is sold to a mineral extractor. If the mine operates for x hours in a day, what is the simplest inequality that represents the number of hours the mine must operate to meet its obligations?

OpenStudy (danjs):

The general idea is the inequality for the mine - \[production \ge obligations\]

OpenStudy (dannimarie):

So what would the exact answer be?

OpenStudy (danjs):

well, the amount produced is 4 tons/hr times X number of hours extracting... production = 4*X tons

OpenStudy (danjs):

The obligations are 4 for the govt, and at least 10 for the other one... so that means they need 14 tons at least \[obligations \ge 14\]

OpenStudy (danjs):

Sorry the production was 2 not 4 tons/hr, so production is 2*X so together it is.. \[2x \ge 14\] \[x \ge 7\] They must operate at least 7 hours each day

OpenStudy (dannimarie):

Thank you!

OpenStudy (dannimarie):

So what would the exact answer be?

OpenStudy (volleysappspbj16):

What does at least mean? DanniMarie